Find the turning point and dertemine whether they are Maxima or minima.


i.y=x2-6x+9


ii.y=x2+5xz+z2


iii.y=x2+z2+v2

1
Expert's answer
2022-02-08T15:10:54-0500

i.

y'=(x2-6x+9'=2x-6

y''=(2x-6)''=2

2x-6=0

x=3

The second derivative does not contain x , so insertion gives 2. 2 is larger than 0. So there is a mimimum at 3.

"y(3)=3^2-6\times3+9=9-18+9=0"

Minimum turning point (3;0).


ii. y(x):

"x^2+5xz=y-z^2"

"x^2+5xz+(\frac{5z}{2})^2=(\frac{5z}{2})^2+y-z^2"


solution set


"x1=\frac{-5z}{2}+(y-z^2+\frac{25z^2}{4})^{0.5}"

"x2=\frac{-5z}{2}-(y-z^2+\frac{25z^2}{4})^{0.5}"


  y(z)

"z^2+5xz=-x^2+y"


"z^2+5xz+(\frac{5x}{2})^2=(\frac{5x}{2})^2+y-x^2"


"x1=\frac{-5x}{2}+(y-x^2+\frac{25x^2}{4})^{0.5}"


"x2=\frac{-5x}{2}-(y-x^2+\frac{25x^2}{4})^{0.5}"


iii.


y(x)


"z^2+v^2=y-x^2"


"z^2=y-x^2-v^2"


"z=+-(y-x^2-v^2)^{0.5}"


y(z)



"z^2+x^2=y-v^2"


"x^2=y-z^2-v^2"


"x=+-(y-z^2-v^2)^{0.5}"


y(v)

"z^2+v^2=y-x^2"


"v^2=y-z^2-x^2"


"v=+-(y-z^2-x^2)^{0.5}"